OpenStudy (anonymous):

h(t)=cot t, [pi/4,3pi/4] find the avg rate of change of the function over the given intervals HELP!!

OpenStudy (anonymous):

average rate is \[\frac{\cot(\frac{3\pi}{4})-\cot(\frac{\pi}{4})}{\frac{3\pi}{4}-\frac{\pi}{4}}\]